Constants |Periodic Table The position of a particular particle as a function of time is given by T (9.60ti + 8.85j-1.00t k)m, where t is in seconds. PartA Determine the particles velocity as a function of time. Express your answer in terms of the unit vectors i, j, and k. m/s u= Submit Request Answer Part B Determine the particles acceleration as a function of time Express your answer in terms of the unit vectors i, j, and k. Submit Request Answer
Answer #1

a)

v = dr / dt = d ( 9.6 t i + 8.85 j - t^2 k) / dt

v = 9.6 i - 2t k

======

b)

a = d v / dt = d ( 9.6 i - 2t k ) / dt

a = - 2 k

=======

do comment in case any doubt, will reply for sure . Goodluck